Two adjacent blocks of mass m1 = 3.0 kg and m2 = 4.0 kg are on a frictionless surface.A force of 6.0 N is applied to m1,and a force of 4.0 N is applied to m2 ;these antiparallel forces squeeze the blocks together.What is the force due to m1 on m2? Due to m2 on m1? What is the acceleration of the system?
1
Expert's answer
2020-10-07T10:23:27-0400

The force due to "m_{1}" on "m_{2}" is "6 N"

the force due to "m_{2}" on "m_{1}" is "4 N"

To get the acceleration, "a" of the system, we find the net force, "F" .

"F = 6N - 4N = 2N"

total mass "= 3 + 4 = 7"

"F = ma"

"a = \frac{F}{m} = \frac{2}{7} = 0.286 m s^{-2}"